# Moving through Cartesian products, coronas and joins in general position

### v1

@@ 2 Bounds for Cartesian products
For $k\geq 2$, let $V(K_{1,k})=\{0\}\cup[k]$, with 00 being the vertex of degree $k$. It follows from [33, Theorem 1] that $\operatorname{gp}(K_{1,k}\operatorname{\,\square\,}K_{1,k})=2k$. We will show that $\operatorname{Mob_{gp}}(K_{1,k}\operatorname{\,\square\,}K_{1,k})=k+1$, so that $\operatorname{gp}(K_{1,k}\operatorname{\,\square\,}K_{1,k})-\operatorname{Mob_{gp}}(K_{1,k}\operatorname{\,\square\,}K_{1,k})=k-1$ and the result follows on setting $k=r+1$.
Let $S$ by any mobile general position set of $K_{1,k}\operatorname{\,\square\,}K_{1,k}$. Any pair of adjacent vertices in the Cartesian product of trees is a maximal general position set [21], so if $|S|>2$ the robots must always occupy an independent set. We can start at the stage that a robot is stationed at $(0,0)$. If there are robots at vertices $(u_{1},v_{1})$ and $(u_{2},v_{2})$ with $u_{1},v_{1},u_{2},v_{2}\in[k]$, $u_{1}\neq u_{2}$ and $v_{1}\neq v_{2}$, then there would be a shortest $(u_{1},v_{1}),(u_{2},v_{2})$-path through $(0,0)$; hence all the remaining robots lie in a set $\{i\}\times[k]$ or $[k]\times\{j\}$ for some $i,j\in[k]$. In either case, we have $|S|\leq k+1$.
Finally, we show that $\operatorname{Mob_{gp}}(K_{1,k}\operatorname{\,\square\,}K_{1,k})\geq k+1$. Consider the set $S=\{(0,0)\}\cup([k]\times\{1\})$. First, move $(0,0)\rightsquigarrow(0,2)$, followed by $(i,1)\rightsquigarrow(i,0)$ for $i\in[2,k]$ and then $(0,2)\rightsquigarrow(1,2)$. By relabelling as necessary, we see that each vertex except for $(0,1)$ may be visited using these sequences. For the remaining vertex, from the initial configuration $S=\{(0,0)\}\cup([k]\times\{1\})$ perform $(0,0)\rightsquigarrow(0,2)$, $(i,1)\rightsquigarrow(i,0)$ for $i\in[2,k]$, and finally $(1,1)\rightsquigarrow(0,1)$. Therefore, $S$ is a mobile general position set. ∎
@@ 3 Cartesian products with paths
By Proposition 2.2, we can assume that $\ell(T)\geq 3$. As remarked in Section 1, $\operatorname{gp}_{\rm o}(T)=\ell(T)$, so by Proposition 2.1(ii) we have $\operatorname{Mob_{gp}}(T\operatorname{\,\square\,}K_{2})\geq\ell(T)$. We label the vertices of $K_{2}$ by $0,1$. Suppose that at least $\ell(T)+1$ robots can traverse $T\operatorname{\,\square\,}K_{2}$ in general position. Since $\operatorname{gp}(T)=\ell(T)$ and each $T$-layer in $T\operatorname{\,\square\,}K_{2}$ is isometric, neither $T$-layer can contain $>\ell(T)$ robots at any stage, that is, each layer $V(T)\times\{0\}$ and $V(T)\times\{1\}$ must contain at least one robot at any time. Trivially we can assume that at least one layer contains two or more robots.
Suppose that the layer $T^{1}$ contains at least two robots. If not all of the robots in $T^{1}$ are already stationed at leaves of $T$, then we may suppose that a robot is at a vertex $(u,1)$, where $u$ is a cut-vertex of $T$. Let $T_{1},\dots,T_{k}$ be the components of $T-u$. As the set of robots is in general position, one of the sets $V(T_{i})\times\{1\}$ must contain the remaining robots of $T^{1}$; without loss of generality, suppose that these other robots are in $V(T_{1})\times\{1\}$. Considering the shortest paths to the robots in $V(T_{1})\times\{1\}$, we see that there cannot be robots positioned at any vertex from ($(\{u\}\cup\bigcup_{i=2}^{k}V(T_{i}))\times\{0\}$. Therefore, the robot at $(u,1)$ can be moved by a sequence of legal moves to $(v,1)$, where $v$ is a leaf of $T$ lying in $T_{2}$. In this fashion, if both layers $T^{i}$, $i\in\{0,1\}$, contained at least two robots, then all of these robots could be moved to vertices corresponding to leaves of $T$. However, if $w$ is any leaf of $T$, then we cannot have robots at both $(w,0)$ and $(w,1)$, as this constitutes a maximal general position set. Therefore, in this case, we conclude that there are at most $\ell(T)$ robots in $T\operatorname{\,\square\,}K_{2}$, a contradiction.
It follows that there must be a layer, say $T^{0}$, that contains just one robot $R$, and $\ell(T)$ robots lie in $T^{1}$. By the preceding argument, $R$ cannot move to the layer $T^{1}$ and no robot in $T^{1}$ can move to $T^{0}$. If $T$ is a path $P_{n}$, then we are left with three robots: two positioned at vertices of degree two from the copy $P_{n}^{1}$, and other one located at a vertex of degree three from the other copy $P_{n}^{0}$. It is readily observed that no robot can visit the vertices of degree two from the copy $P_{n}^{0}$. Hence, we may assume $T$ is not a path. Let $z$ be any vertex of $T$ with degree at least three. No robot in $T^{1}$ can visit $(z,1)$ without creating three in a line within $T^{1}$, and robot $R$ cannot leave $T^{0}$ to visit $(z,1)$, a contradiction. We conclude that $T\operatorname{\,\square\,}K_{2}$ can hold at most $\ell(T)$ robots. ∎
@@
If $n\geq m\geq 3$, then $\operatorname{Mob_{gp}}(P_{n}\operatorname{\,\square\,}P_{m})=3$.
@@
Let $V(P_{k})=[k]$, so that $V(P_{n}\operatorname{\,\square\,}P_{m})=[n]\times[m]$. Consider an arbitrary general position set $S$ of $P_{n}\operatorname{\,\square\,}P_{m}$ with $|S|=4$. Then from the proof of [17, Theorem 2.1] we deduce that none of the vertices $(1,1)$, $(1,m)$, $(n,1)$ and $(n,m)$ belongs to $S$. Hence, no sequence of legal moves for any configuration of four robots in general position in $P_{n}\operatorname{\,\square\,}P_{m}$ can visit any of the vertices $(1,1)$, $(1,m)$, $(n,1)$ and $(n,m)$. Thus, $\operatorname{Mob_{gp}}(P_{n}\operatorname{\,\square\,}P_{m})\leq 3$.
@@
  • $(1,1)$ moves to all the vertices from $[1]\times[m-1]$ and returns back to $(1,1)$.
@@
  • $(2,m)$ moves to all the vertices from $[2,n-1]\times[2,m]$ and returns back to $(2,m)$.
  • $(n,1)\rightsquigarrow(n,2)$. After this, $(1,1)$ moves to vertices $(2,1),\dots,(n-1,1)$. Notice that at this point the robots are at vertices $(2,m)$, $(n-1,1)$ and $(n,2)$. Moreover, by this stage, all the vertices apart from $(1,m)$ and $(n,m)$ have already been visited.
@@
Notice that all these moves are legal, which demonstrates that $\operatorname{Mob_{gp}}(P_{n}\operatorname{\,\square\,}P_{m})\geq 3$ and hence $\operatorname{Mob_{gp}}(P_{n}\operatorname{\,\square\,}P_{m})=3$ when $n\geq m\geq 3$. ∎
@@
We first recall from [26, Corollary 3.2] that $\operatorname{gp}(P_{\infty}\operatorname{\,\square\,}P_{\infty})=4$. Hence, it remains to show that $\operatorname{Mob_{gp}}(P_{\infty}\operatorname{\,\square\,}P_{\infty})\geq 4$. To do so, let $V(P_{\infty})=\mathbb{Z}$ and let $(i,j)\in V(P_{\infty}\operatorname{\,\square\,}P_{\infty})$. We claim that the set $S=\{(i-1,j),(i+1,j),(i,j-1),(i,j+1)\}$ is a mobile general position set of $P_{\infty}\operatorname{\,\square\,}P_{\infty}$. Observe that the sequence of moves $(i+1,j)\rightsquigarrow(i+2,j)$, $(i,j+1)\rightsquigarrow(i+1,j+1)$, $(i,j-1)\rightsquigarrow(i+1,j-1)$ and $(i-1,j)\rightsquigarrow(i,j)$ is a sequence of legal moves from robots positioned at the set $S$. This sequence moves the four robots one coordinate to the right. We can analogously move the four robots in each of the remaining three directions in the infinite grid. In this way, every vertex of the infinite grid is eventually occupied by some robot. Thus, the conclusion follows. ∎
@@
We assume for the remainder of the proof that $n\geq 5$. It follows from [17, Theorem 3.2] that $\operatorname{gp}(C_{n}\operatorname{\,\square\,}K_{2})=4$. Let $V(C_{n})=\mathbb{Z}_{n}$ and $V(K_{2})=[2]$. We separate the argument into two cases.
@@
Consider a set of three robots located at $S=\{(v_{0},1),(v_{\lceil n/2\rceil},1),(v_{1},2),(v_{\lceil n/2\rceil+1},2)\}\,.$ Then $S$ is a general position set. Moreover, consider the following sequence of four moves for the robots:
@@
Suppose now that the robots are positioned at $S=\{(v_{0},1),(v_{n/2},1),(v_{1},2),(v_{n/2+1},2)\}\,.$ Then $S$ is a general position set. Moreover, consider the following sequence of moves:
@@
If $H$ is a connected graph with girth at least $2r$ and radius at least $r-1$, then for any graph $G$ it holds that
@@
Let $Q$ be any path $v_{1},v_{2},\dots,v_{r}$ of length $r-1$ in $H$. The subgraph of $G\operatorname{\,\square\,}H$ induced by $V(G)\times V(Q)$ is isomorphic to $G\operatorname{\,\square\,}P_{r}$, and as the girth of $H$ is at least $2r$, the subgraph is isometric. Thus, $\operatorname{Mob_{gp}}(G\operatorname{\,\square\,}P_{r})$ robots can traverse the vertices of $V(G)\times V(Q)$ in general position without leaving the subgraph. Now, let $v_{r+1}\in N_{H}(v_{r})\setminus\{v_{r-1}\}$. Suppose that the robots have visited all the vertices of $V(G)\times V(Q)$ by a sequence of legal moves. Next, move all robots in $V(G)\times\{v_{r}\}$ to $V(G)\times\{v_{r+1}\}$ in turn by legal moves of the form $(u,v_{r})\rightsquigarrow(u,v_{r+1})$, where $u\in V(G)$. Then repeat this process to move the robots in $V(G)\times\{v_{j}\}$ to $V(G)\times\{v_{j+1}\}$ for $j=r-1,r-2,\dots,1$. As $H$ has girth at least $2r$, the robots remain in general position throughout this process. As $H$ is connected and the radius of $H$ is at least $r-1$, any layer $G^{h}$ can be visited in this way. ∎
@@
Clearly, no three robots from $S$ can lie in the same $P_{s}$-layer. Suppose that two robots $R_{1}$ and $R_{2}$ in $S$ lie in the same $P_{s}$-layer of $C_{4}\operatorname{\,\square\,}P_{s}$; without loss of generality, $R_{1}$ and $R_{2}$ are stationed at vertices $(0,i)$ and $(0,j)$ respectively, where $1\leq i<j\leq s$. There is a shortest path from any vertex $(u,v)$ with $v\in[i]$ to $R_{2}$ through $R_{1}$, and likewise for vertices with second coordinate at least $j$. Hence, there are no robots on $\mathbb{Z}_{4}\times\left([1,i]\cup[j,s]\right)$ apart from $R_{1}$ and $R_{2}$. Call the other two robots $R_{3}$ and $R_{4}$. By the above analysis, $R_{3}$ and $R_{4}$ cannot have the same first coordinate. Hence we can assume that robot $R_{3}$ lies at $(1,k)$, where $i<k<j$. Any vertex in ($\left(\{1,2\}\times[i+1,j-1]\right)\setminus\{(1,k)\}$ has a shortest path to either $R_{1}$ or $R_{2}$ through $R_{3}$, so $R_{4}$ must be at a vertex $(3,l)$, where $i<l<j$. However, in this configuration each robot is only free to move within its $P_{s}$-layer, and so no robot can visit any vertex in ${}^{2}P_{s}$, a contradiction. Consequently, there is one robot on each $P_{s}$-layer, and none of the robots can move to another $P_{s}$-layer.
Observe that any pair of adjacent vertices constitutes a maximal general position set of $C_{4}\operatorname{\,\square\,}P_{s}$, so the robots must at all times occupy an independent set. Therefore, if we suppose that $R_{1}$ is the robot located at the vertex with smallest second coordinate in the initial configuration, say at vertex $(0,i)$, then no robot can ever move to a position with second coordinate smaller than that of $R_{1}$, and hence the vertices in $\{1,2,3\}\times\{1\}$ cannot be visited by legal moves. Thus, $\operatorname{Mob_{gp}}(C_{4}\operatorname{\,\square\,}P_{s})\leq 3$.
@@
Theorem 3.4 for prisms along with Lemma 3.5 implies that $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}P_{s})\geq 4$ for $r\geq 5$. Combined with the upper bound involving $\operatorname{gp}(C_{r}\operatorname{\,\square\,}P_{s})$ we see that $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}P_{s})=4$ for $r\in\{5,6,8\}$ and $s\geq 2$, or for $r\geq 5$ and $s\leq 4$. For $r=7$ or $r\geq 9$ and $s\geq 5$ the mobile general position number of $C_{r}\operatorname{\,\square\,}P_{s}$ must be either four or five.
@@
If $r=9$ or $r\geq 11$ and $s\geq 5$ are integers, then $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}P_{s})=5$.
@@
It is known from [17, Theorem 3.2] that $\operatorname{gp}(C_{r}\operatorname{\,\square\,}P_{s})=5$ for $r=9$ or $r\geq 11$ and $s\geq 5$. It only remains to show that $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}P_{s})\geq 5$. We show that the result is true for $C_{r}\operatorname{\,\square\,}P_{5}$, and the full claim then follows for larger values of $s$ by Lemma 3.5. Set $V(C_{r})=\mathbb{Z}_{r}$ and $V(P_{5})=[5]$.
@@
For $r=9$ or $r>10$ and $s\geq 10$, $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}C_{s})\geq 5$.
It is shown in [20] that if $r,s\geq 7$ and $r$ and $s$ do not both lie in $\{8,10,12\}$, then $\operatorname{gp}(C_{r}\operatorname{\,\square\,}C_{s})=7$. Computer search shows that the torus $C_{9}\operatorname{\,\square\,}C_{8}$ has mobile general position number seven.

### v2

@@ 2 Bounds for Cartesian products
For $k\geq 2$, let $V(K_{1,k})=\{0\}\cup[k]$, with 0 being the vertex of degree $k$. It follows from [31, Theorem 1] that $\operatorname{gp}(K_{1,k}\operatorname{\,\square\,}K_{1,k})=2k$. We will show that $\operatorname{Mob_{gp}}(K_{1,k}\operatorname{\,\square\,}K_{1,k})=k+1$, so that $\operatorname{gp}(K_{1,k}\operatorname{\,\square\,}K_{1,k})-\operatorname{Mob_{gp}}(K_{1,k}\operatorname{\,\square\,}K_{1,k})=k-1$ and the result follows on setting $k=r+1$.
Let $S$ be any mobile general position set of $K_{1,k}\operatorname{\,\square\,}K_{1,k}$. Any pair of adjacent vertices in the Cartesian product of trees is a maximal general position set [21], so if $|S|>2$ the robots must always occupy an independent set. We can start at the stage that a robot is stationed at $(0,0)$. If there are robots at vertices $(u_{1},v_{1})$ and $(u_{2},v_{2})$ with $u_{1},v_{1},u_{2},v_{2}\in[k]$, $u_{1}\neq u_{2}$ and $v_{1}\neq v_{2}$, then there would be a shortest $(u_{1},v_{1}),(u_{2},v_{2})$-path through $(0,0)$; hence all the remaining robots lie in a set $\{i\}\times[k]$ or $k]\times\{j\}$ for some $i,j\in[k]$. In either case, we have $|S|\leq k+1$.
Finally, we show that $\operatorname{Mob_{gp}}(K_{1,k}\operatorname{\,\square\,}K_{1,k})\geq k+1$. Consider the set $S=\{(0,0)\}\cup([k]\times\{1\})$. First, move $(0,0)\rightsquigarrow(0,2)$, followed by $(i,1)\rightsquigarrow(i,0)$ for $i\in[2,k]$ and then $(0,2)\rightsquigarrow(1,2)$. By relabelling as necessary, we see that each vertex except for $(0,1)$ may be visited using such a sequence. For the remaining vertex, from the initial configuration $S=\{(0,0)\}\cup([k]\times\{1\})$ perform $(0,0)\rightsquigarrow(0,2)$, $(i,1)\rightsquigarrow(i,0)$ for $i\in[2,k]$, and finally $(1,1)\rightsquigarrow(0,1)$. Therefore, $S$ is a mobile general position set. ∎
@@ 3 Cartesian products with paths
By Proposition 2.2, we can assume that $\ell(T)\geq 3$. As remarked in Section 1, $\operatorname{gp}_{\rm o}(T)=\ell(T)$, so by Proposition 2.1(ii) we have $\operatorname{Mob_{gp}}(T\operatorname{\,\square\,}K_{2})\geq\ell(T)$. We label the vertices of $K_{2}$ by $0,1$. Suppose that at least $\ell(T)+1$ robots can traverse $T\operatorname{\,\square\,}K_{2}$ in general position. Since $\operatorname{gp}(T)=\ell(T)$ and each $T$-layer in $T\operatorname{\,\square\,}K_{2}$ is a convex subgraph, neither $T$-layer can contain $\ell(T)$ robots at any stage, that is, each layer $V(T)\times\{0\}$ and $V(T)\times\{1\}$ must contain at least one robot at any time. Trivially we can assume that at least one layer contains two or more robots.
Suppose that the layer $T^{1}$ contains at least two robots. If not all of the robots in $T^{1}$ are already stationed at leaves of $T$, then we may suppose that a robot is at a vertex $(u,1)$, where $u$ is a cut-vertex of $T$. Let $T_{1},\dots,T_{k}$ be the components of TuT-u. As the set of robots is in general position, one of the sets $V(T_{i})\times\{1\}$ must contain the remaining robots of $T^{1}$; without loss of generality, suppose that these other robots are in $V(T_{1})\times\{1\}$. Considering the shortest paths to the robots in $V(T_{1})\times\{1\}$, we see that there cannot be robots positioned at any vertex from $(\{u\}\cup\bigcup_{i=2}^{k}V(T_{i}))\times\{0\}$. Therefore, the robot at $(u,1)$ can be moved by a sequence of legal moves to $(v,1)$, where $v$ is a leaf of $T$ lying in $T_{2}$. In this fashion, if both layers $T^{i}$, $i\in\{0,1\}$, contained at least two robots, then all of these robots could be moved to vertices corresponding to leaves of $T$. However, if $w$ is any leaf of $T$, then we cannot have robots at both $(w,0)$ and $(w,1)$, as this constitutes a maximal general position set of $T\operatorname{\,\square\,}K_{2}$. Therefore, in this case, we conclude that there are at most $\ell(T)$ robots in $T\operatorname{\,\square\,}K_{2}$, a contradiction.
It follows that there must be a layer, say $T^{0}$, that contains just one robot $R$, and $\ell(T)$ robots lie in $T^{1}$, which we can assume to start at the leaves of $T^{1}$. By the preceding argument, $R$ cannot move to the layer $T^{1}$ and no robot in $T^{1}$ can move to $T^{0}$. If $T$ is a path $P_{n}$, then we are left with three robots: two positioned at vertices corresponding to leaves of the layer $P_{n}^{1}$, and the third robot located at an internal vertex of the path layer $P_{n}^{0}$. It is now readily observed that no robot can visit the vertices corresponding to leaves of $P_{n}^{0}$. Hence, we may assume that $T$ is not a path. Let $z$ be any vertex of $T$ with degree at least three. No robot in $T^{1}$ can visit $(z,1)$ without creating three-in-a-line within $T^{1}$, and robot $R$ cannot leave $T^{0}$ to visit $(z,1)$, a contradiction. We conclude that $T\operatorname{\,\square\,}K_{2}$ can hold at most $\ell(T)$ robots. ∎
@@
If $n,m\geq 3$, then $\operatorname{Mob_{gp}}(P_{n}\operatorname{\,\square\,}P_{m})=3$.
@@
Let $V(P_{k})=[k]$, so that $V(P_{n}\operatorname{\,\square\,}P_{m})=[n]\times[m]$. Consider an arbitrary general position set $S$ of $P_{n}\operatorname{\,\square\,}P_{m}$ with $|S|=4$. Then from the proof of [18, Theorem 2.1] we deduce that none of the corner vertices $(1,1)$, $(1,m)$, $(n,1)$ and $(n,m)$ belongs to $S$. Hence, no sequence of legal moves for any configuration of four robots in general position in $P_{n}\operatorname{\,\square\,}P_{m}$ can visit any of the vertices $(1,1)$, $(1,m)$, $(n,1)$ and $(n,m)$. Thus, $\operatorname{Mob_{gp}}(P_{n}\operatorname{\,\square\,}P_{m})\leq 3$.
@@
  • $(1,1)$ moves to all the vertices from $\{1\}\times[m-1]$ and returns back to $(1,1)$.
@@
  • $(2,m)$ moves to all the vertices from $2,n-1]\times[2,m]$ and returns back to $(2,m)$.
  • $(n,1)\rightsquigarrow(n,2)$. After this, $(1,1)$ moves to vertices $(2,1),\dots,(n-1,1)$. Notice that at this point the robots are at vertices $(2,m)$, (n1,1)(n-1,1) and $(n,2)$. Moreover, by this stage, all the vertices apart from $(1,m)$ and $(n,m)$ have already been visited.
@@
Notice that all these moves are legal, which demonstrates that $\operatorname{Mob_{gp}}(P_{n}\operatorname{\,\square\,}P_{m})\geq 3$ and hence $\operatorname{Mob_{gp}}(P_{n}\operatorname{\,\square\,}P_{m})=3$ when $n,m\geq 3$. ∎
@@
We first recall from [26, Corollary 3.2] that $\operatorname{gp}(P_{\infty}\operatorname{\,\square\,}P_{\infty})=4$. Hence, it remains to show that $\operatorname{Mob_{gp}}(P_{\infty}\operatorname{\,\square\,}P_{\infty})\geq 4$. To do so, set $V(P_{\infty})=\mathbb{Z}$ and let $(i,j)\in V(P_{\infty}\operatorname{\,\square\,}P_{\infty})$. We claim that the set $N(i,j)=\{(i-1,j),(i+1,j),(i,j-1),(i,j+1)\}$ is a mobile general position set of $P_{\infty}\operatorname{\,\square\,}P_{\infty}$. First, from the proof of [26, Corollary 3.2], we know that any such set $N(i,j)$ is in general position. Next, observe that the sequence of moves $(i+1,j)\rightsquigarrow(i+2,j)$, $(i,j+1)\rightsquigarrow(i+1,j+1)$, $(i,j-1)\rightsquigarrow(i+1,j-1)$ and $(i-1,j)\rightsquigarrow(i,j)$ is a sequence of legal moves from robots positioned at the set $N(i,j)$. This sequence moves the four robots one coordinate to the right, leaving robots at $N(i+1,j)$. Fig. 4 shows a gp-set of the grid $P_{\infty}\operatorname{\,\square\,}P_{\infty}$, and one round of moves as just described. The order of the moves is shown by the numeric order in the figure. We can analogously move the four robots in each of the remaining three directions in the infinite grid. In this way, every vertex of the infinite grid is eventually occupied by some robot. Thus, the conclusion follows. ∎
1324
Figure 4: The legal moves in the infinite grid.
@@
We assume for the remainder of the proof that $n\geq 5$. It follows from [18, Theorem 3.2] that $\operatorname{gp}(C_{n}\operatorname{\,\square\,}K_{2})=4$. Set $V(C_{n})=\{v_{i}:i\in\mathbb{Z}_{n}\}$ and $V(K_{2})=[2]$. We separate the argument into two cases.
@@
Consider a set of four robots located at $S=\{(v_{0},1),(v_{\lceil n/2\rceil},1),(v_{1},2),(v_{\lceil n/2\rceil+1},2)\}\,.$ Then $S$ is a general position set. Moreover, consider the following sequence of four moves for the robots:
@@
Suppose now that the robots are positioned at $S=\{(v_{0},1),(v_{n/2},1),(v_{1},2),(v_{n/2+1},2)\}\,.$ Again $S$ is a general position set. Moreover, consider the following sequence of moves:
@@
If $H$ is a connected graph with girth at least $2r$ and radius at least r1r-1, then for any graph $G$ it holds that
@@
Let $Q$ be any path $v_{1},v_{2},\dots,v_{r}$ of length r1r-1 in $H$. The subgraph of $G\operatorname{\,\square\,}H$ induced by $V(G)\times V(Q)$ is isomorphic to $G\operatorname{\,\square\,}P_{r}$, and as the girth of $H$ is at least $2r$, the subgraph is convex. Thus, $\operatorname{Mob_{gp}}(G\operatorname{\,\square\,}P_{r})$ robots can traverse the vertices of $V(G)\times V(Q)$ in general position without leaving the subgraph. Now, let $v_{r+1}\in N_{H}(v_{r})\setminus\{v_{r-1}\}$. Suppose that the robots have visited all the vertices of $V(G)\times V(Q)$ by a sequence of legal moves. Next, move all robots in $V(G)\times\{v_{r}\}$ to $V(G)\times\{v_{r+1}\}$ in turn by legal moves of the form $(u,v_{r})\rightsquigarrow(u,v_{r+1})$, where $u\in V(G)$. Then repeat this process to move the robots in $V(G)\times\{v_{j}\}$ to $V(G)\times\{v_{j+1}\}$ for $j=r-1,r-2,\dots,1$. As $H$ has girth at least $2r$, the robots remain in general position throughout this process. As $H$ is connected and the radius of $H$ is at least r1r-1, any layer $G^{h}$ can be visited in this way. ∎
@@
Clearly, no three robots from $S$ can lie in the same $P_{s}$-layer. Suppose that two robots $R_{1}$ and $R_{2}$ in $S$ lie in the same $P_{s}$-layer of $C_{4}\operatorname{\,\square\,}P_{s}$; without loss of generality, $R_{1}$ and $R_{2}$ are stationed at vertices $(0,i)$ and $(0,j)$ respectively, where $1\leq i<j\leq s$. There is a shortest path from any vertex $(u,v)$ with $v\in[i]$ to $R_{2}$ through $R_{1}$, and likewise for vertices with second coordinate at least $j$. Hence, there are no robots on $\mathbb{Z}_{4}\times\left([1,i]\cup[j,s]\right)$ apart from $R_{1}$ and $R_{2}$. Call the other two robots $R_{3}$ and $R_{4}$. By the above analysis, $R_{3}$ and $R_{4}$ cannot have the same first coordinate. Hence we can assume that robot $R_{3}$ lies at $(1,k)$, where $i<k<j$. Any vertex in $\left(\{1,2\}\times[i+1,j-1]\right)\setminus\{(1,k)\}$ has a shortest path to either $R_{1}$ or $R_{2}$ through $R_{3}$, so $R_{4}$ must be at a vertex $(3,l)$, where $i<l<j$. However, in this configuration each robot is only free to move within its $P_{s}$-layer, and so no robot can visit any vertex in $^{2}P_{s}$, a contradiction. Consequently, there is one robot on each $P_{s}$-layer, and none of the robots can move to another $P_{s}$-layer.
Observe that any pair of adjacent vertices constitutes a maximal general position set of $C_{4}\operatorname{\,\square\,}P_{s}$, so the robots must at all times occupy an independent set. Therefore, if we suppose that $R_{1}$ is the robot located at the vertex with smallest second coordinate in the initial configuration, say at vertex $(0,i)$, then no robot in the layers $^{1}P_{s}$ or $^{3}P_{s}$ can ever move to a position with second coordinate smaller than that of $R_{1}$, and hence the vertices in $\{1,3\}\times\{1\}$ cannot be visited by legal moves. Thus, $\operatorname{Mob_{gp}}(C_{4}\operatorname{\,\square\,}P_{s})\leq 3$.
@@
Theorem 3.4 for prisms along with Lemma 3.5 implies that $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}P_{s})\geq 4$ for $r\geq 5,s\geq 2$. Combined with the upper bound involving $\operatorname{gp}(C_{r}\operatorname{\,\square\,}P_{s})$ we see that $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}P_{s})=4$ for $r\in\{5,6,8\}$ and $s\geq 2$, or for $r\geq 5$ and $s\leq 4$. In the cases $r=7$ or $r\geq 9$ and $s\geq 5$ the mobile general position number of $C_{r}\operatorname{\,\square\,}P_{s}$ must be either four or five.
@@
If $r=9$ or $r\geq 11$, and $s\geq 5$, then $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}P_{s})=5$.
@@
It is known from [18, Theorem 3.2] that $\operatorname{gp}(C_{r}\operatorname{\,\square\,}P_{s})=5$ for $r=9$ or $r\geq 11$, and $s\geq 5$. It only remains to show that $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}P_{s})\geq 5$. We show that the result is true for $C_{r}\operatorname{\,\square\,}P_{5}$, and the full claim then follows for larger values of $s$ by Lemma 3.5. Set $V(C_{r})=\mathbb{Z}_{r}$ and $V(P_{5})=[5]$.
@@
For $r=9$ or $r>10$, and $s\geq 10$, $\operatorname{Mob_{gp}}(C_{r}\operatorname{\,\square\,}C_{s})\geq 5$.
It is shown in [20] that if $r,s\geq 7$ and $r$ and $s$ do not both lie in $\{8,10,12\}$, then $\operatorname{gp}(C_{r}\operatorname{\,\square\,}C_{s})=7$. Computer search shows that the torus $C_{9}\operatorname{\,\square\,}C_{8}$ has mobile general position number seven, so this upper bound can be achieved.